OpenStudy (anonymous):

Find the circumference and area of a circle with a diameter of 14 inches. Leave your answers in terms of pi.

OpenStudy (anonymous):

C = 14π; A = 49π C = 7π; A = 28π C = 14π; A = 28π C = 7π; A = 49π

OpenStudy (anonymous):

the radius is 1/2 the diameter, so the radius is 7. The area of a circle is pi(r^2) = 49*pi The circumference of a circle is 2(pi)r = (2)(7)(pi) = 14(pi) first choice is the answer

OpenStudy (anonymous):

what about this one Given a soda can with a volume of 15 and a diameter of 2, what is the volume of a cone that fits perfectly inside the soda can? (Hint: only enter numerals in the answer blank).

OpenStudy (anonymous):

sorry...that I don't know....not real good at doing those

OpenStudy (anonymous):

ok thank you anyway.